OpenStudy (anonymous):

Find 2p ^2 + 3p - 4 less - 2p ^2 - 3p + 4.

OpenStudy (marihelenh):

Same thing but now with subtraction. You will take 2p^2-(-2p^2) and 3p-(-3p) and -4-4.

OpenStudy (marihelenh):

Subtracting a negative number is the same as adding it, so you can 2p^2+2p^2 and 3p+3p.

OpenStudy (mathstudent55):

\((2p^2 + 3p - 4) - (-2p^2 - 3p + 4) =\) You need to subtract the polynomials above. The first step is to rewrite the first polynomial without parentheses since those parentheses are not needed. \(= 2p^2 + 3p - 4 - (-2p^2 - 3p + 4) \) The next step is to deal with the parentheses around the second polynomial. We need to get rid of those parentheses.

OpenStudy (marihelenh):

When you add them, you should get 4p^2+6p-8.

OpenStudy (anonymous):

so 4p^2 and 6p and -8?

OpenStudy (anonymous):

thanks for the help

OpenStudy (mathstudent55):

Since there is a negative sign to the left of the parentheses, in order to get rid of the parentheses, you distribute the negative sign to each each term inside the parentheses. What ends up happening is that every sign inside the parentheses changes. \(= 2p^2 + 3p - 4 - (-2p^2 - 3p + 4)\) \(= 2p^2 + 3p - 4 + 2p^2 + 3p - 4\) Now that the parentheses are taken care of, you just combine like terms like we did before. \(= 2p^2 + 2p^2 + 3p + 3p - 4 - 4\) \(= 4p^2 + 6p + - 8\)